# Fine Spectrum from Crude Analytic Bootstrap

### v1

@@ 4.5 Large charge
A sector that is relatively easy to study is that of the operators of largest $O(N-1)$ charge at a given dimension close to $J$. There is one operator with charge $J$ and one with $J-1$ and then more with $J-2$, listed in Table 5. This is the pattern we see also for $\Delta\sim 2,3,4$ in the subsections above, and as in those cases, there are more operators of charge $J-2$, which in terms of fundamental fields are obtained by acting on $\phi_{a_{1}}\cdots\phi_{a_{J-2}}$ with either $\partial_{\tau}^{2}$ or $\partial_{\mu}\partial^{\mu}$ before subtracting the trace terms.
Name $O(N-1)$ $O(d-1)$ Dimension Fields
@@
We label the operators with a superscript $J$ which is their approximate dimension, not their $O(N-1)$ quantum number. The charge can be seen in the number of indices on the operators in Table 5, though we sometimes avoid the multiple indices by using a null polarisation vector $\theta^{a}$ and contract to
@@ 5 Integral identities
[…]
@@
The critical coupling $h_{*}$ was computed to this order in [3, 1] and the $21-\varepsilon/$ 2 comes from the renormalisation of $t$ in the MS scheme [4] and the expansion of the free propagator at $d=4-\varepsilon$.
@@
[…]
@@
The same integral identity (5.3) holds, where the $\theta$ derivative on the right-hand side now gives a factor of $J-1$. At order $\varepsilon^{0}$ the only contribution is from the contact terms and amounts to
@@ 6 Conclusion
All the results derived are perturbative in $\varepsilon$ and rely on Feynman diagrams. It is worth comparing the method used here to direct Feynman diagram calculations of all operators presented in Section 4. To implement our algorithm (with extensive redundancies), we computed around seventy four-point functions. This is a large number, but the computations are extremely simple and in fact algorithmic (though we did not fully automate them).
The most subtle ingredient required in the calculation is the renormalisation factors of the external operators, of which we had only eight (see Tables 1, 2), and only in one case (of $s_{+}$, $s_{-}$) was operator mixing important. So only eleven two-point functions had to be calculated. Additionally, as mentioned in Section 4, two operators of dimension close to three do not appear in any OPE channel of lower dimensional operators and also their dimensions had to be calculated directly along with their renormalisation factors. For comparison, to compute the mixing matrices of the seven nearly degenerate scalars $s^{\prime\prime}_{M}$ in the first row of Table 4 would require 28 two-point functions.
The computational complexity does not seem significantly different, with our method offering the advantage of multiple independent consistency checks and providing some information about structure constants without extra effort. The clear disadvantage is that it provides the spectrum, but not the map between operators and their expression in terms of basic fields. This still demands the evaluation of the renormalisation factors and so requires them to be operators in a Feynman diagram calculation.
This method can easily be applied to other defects in the same model [28, 29, 30, 31, 32, 33, 34] or to other line defects in different models [35]. Likewise, it can be applied to the large $N$ expansion of this and other models.

### v2

@@ 4.5 Large charge
A sector that is relatively easy to study is that of the operators of largest O(N1)O(N-1) charge at a given dimension close to $J$. There is one operator with charge $J$ and one with J1J-1 and then more with J2J-2, listed in Table 5. This is the pattern we see also for $\Delta\sim 2,3,4$ in the subsections above, and as in those cases, there are more operators of charge J2J-2, which in terms of fundamental fields are obtained by acting on $\phi_{a_{1}}\cdots\phi_{a_{J-2}}$ with either $\partial_{\tau}^{2}$ or $\partial_{\mu}\partial^{\mu}$ before subtracting the trace terms.
Name O(N1)O(N-1) O(d1)O(d-1) Dimension Fields
@@
We label the operators with a superscript $J$ which is their approximate dimension, not their O(N1)O(N-1) quantum number. The charge can be seen in the number of indices on the operators in Table 5, though we sometimes avoid the multiple indices by using a null polarisation vector $\theta^{a}$ and contract to
@@ 5 Integral identities
${}2\int d\chi\log|\chi|$ $\langle t_{a}(1)t_{b}(\chi){\mathcal{T}}^{J}(0,\theta){\mathcal{T}}^{J}(\infty,\theta^{\prime})\rangle_{c}=(\theta_{a}\partial_{\theta^{b}}-\theta_{b}\partial_{\theta^{a}})\langle{\mathcal{T}}^{J}(0,\theta){\mathcal{T}}^{J}(\infty,\theta^{\prime})\rangle\,.$ (5.1)
@@
The critical coupling $h_{*}$ was computed to this order in [3, 1] and the $1-\varepsilon/2$ comes from the renormalisation of $t$ in the MS scheme [4] and the expansion of the free propagator at $d=4-\varepsilon$.
@@
$\int d\tau_{1}\,d\tau_{2}\,(\varphi(\tau_{1})-\varphi(\tau_{2}))\langle t_{a}(\tau_{1})t_{b}(\tau_{2}){\mathcal{T}}^{J}(\tau_{3},\theta){\mathcal{T}}^{J}(\tau_{4},\theta^{\prime})\rangle_{c}=(\theta_{a}\partial_{\theta^{b}}-\theta_{b}\partial_{\theta^{a}})\langle{\mathcal{T}}^{J}(\tau_{3},\theta){\mathcal{T}}^{J}(\tau_{4},\theta^{\prime})\rangle\,.$ (5.3)
@@
The same integral identity (5.3) holds, where the $\theta$ derivative on the right-hand side now gives a factor of J1J-1. At order $\varepsilon^{0}$ the only contribution is from the contact terms and amounts to
@@ 6 Conclusion
All the results derived are perturbative in $\varepsilon$ and rely on Feynman diagrams. It is worth comparing the method used here to direct evaluation of anomalous dimensions and structure constants of all operators presented in Section 4 via renormalising and diagonalising two-point and three-point functions. To implement our algorithm (with extensive redundancies), we computed around seventy four-point functions. This is a large number, but the computations are extremely simple and in fact algorithmic (though we did not fully automate them).
The most subtle ingredient required in the calculation is the renormalisation factors of the external operators, of which we had only eight (see Tables 1, 2), and only in one case (of $_{+}, ss_{-}$) was operator mixing important. So only eleven two-point functions had to be calculated. Additionally, as mentioned in Section 4, two operators of dimension close to three do not appear in any OPE channel of lower dimensional operators and also their dimensions had to be calculated directly along with their renormalisation factors. For comparison, to compute the mixing matrices of the seven nearly degenerate scalars $s^{\prime\prime}_{M}$ in the first row of Table 4 would require 28 two-point functions.
The computational complexity does not seem significantly different, with our method offering the advantage of multiple independent consistency checks and providing some information about structure constants without extra effort. The clear disadvantage is that it provides the spectrum, but not the map between operators and their expression in terms of basic fields. This still demands the evaluation of the renormalisation factors and so requires them to be operators in a Feynman diagram calculation. Higher-point functions could substitute the need to perform such explicit renormalisations of new operators. For example, six-point functions of dimension-one operators may help unmix dimension-three operators, similarly to how we were able to unmix dimension-two operators using four-point functions of dimension-one operators.
This method can easily be applied to other defects in the same model [28, 29, 30, 31, 32, 33, 34] or to other line defects in different models [35]. Likewise, it can be applied to the large $N$ expansion of this and other models. In another context, it may prove useful in extracting data in long-range models as discussed in recent literature [36].